Storico Co. just paid a dividend of $2.05 per share. The company will increase its dividend by 24 percent next year and will then reduce its dividend growth rate by 6 percentage points per year until it reaches the industry average of 6 percent dividend growth, after which the company will keep a constant growth rate forever. If the stock price is $57.63, what required return must investors be demanding on Storico stock? (Hint: Set up the valuation formula with all the relevant cash flows, and use trial and error to find the unknown rate of return.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Solution :

It is given that dividend in year 0 = 2.05

Growth rate for next year will be 24% and it will reduce by 6% every year. It means that growth in year 1 , 2, 3, 4 will be 24% , 18%, 12% and 6% respectively.

Let's find the dividend and terminal value

D1 = D0 * ( 1+24% ) = 2.05 *1.24 = 2.54

D2 = D1* (1+18% ) = 2.54 * 1.18= 3.00

D3 = D2* ( 1+ 12%) = 3.00 * 1.12 = 3.36

D4 = D3 * ( 1+6%) = 3.36 * 1.06 = 3.56

Terminal value = D4 * ( 1+ growth rate ) / ( required rate of return - growth rate)

Share price = D1 / ( 1+r)^1 + D2 /( 1+r)^2 + D3 / (1+r)^3 + D4/ (1+r)^4 + TV / ( 1+r)^4

In Excel we can use goal seek function to find the cost of equity or required rate of return.

Terminal Value Year Dividend Discounted Value 3.36 =F3*(1.06)/(B7-0.06) 2.54 3.00 2.29 2.43 =C3/(1+$B$7)^C2 =D3/(1+$B$7)^D2 3

The required rate of return is 11.14%